# RAG-Reward: Optimizing RAG with Reward Modeling and RLHF

### v1

@@ 5 Limitation of Existing Reward Models
In this section, we evaluate several existing reward models in our test set. We select models from RewardBench (Lambert et al., 2024) known for their strong performance in assessing aspects such as helpfulness, safety, and reasoning. We examine their performance on diverse RAG scenarios using our curated test set, and demonstrate their limitation of the evaluation on RAG domains.
The experiment results are shown in Table 5. While many of the listed reward models achieve accuracy higher than 90% in evaluating chat, safety, and reasoning tasks, their overall accuracy in RAG scenarios is less than 80%. This underscores a significant gap between mainstream reward models and the unique requirements of RAG tasks. We also observe inconsistent performance across tasks for the reward models, with better results on Data-to-Text tasks (e.g., the Yelp dataset) compared to the other two types, suggesting that current reward models are not uniformly capable across different RAG scenarios. Interestingly, several reward models that achieve state-of-the-art (SOTA) performance in reasoning and safety evaluations, as shown on the leaderboard [^2] https://huggingface.co/spaces/allenai/reward-bench, do not perform well on RAG tasks. For example, models like URM-LLaMa-3.1-8B (Lou et al., 2024) and Skywork-Reward-Llama-3.1-8B-v0.2 (Liu et al., 2024) underperform in the RAG domain.
In contrast, UltraRM-13b (Cui et al., 2024), which performs suboptimally in reasoning and safety evaluations but excels in assessing helpfulness and instruction-following, achieves near-top accuracy on RAG tasks. This suggests that reward models trained primarily on reasoning tasks may not generalize effectively to evaluating RAG-specific generations. Domain-specific training data are therefore essential to address this gap and improve RAG performance evaluation.
@@ 6.1 Reward Modeling
During the test stage, each test sample contains a chosen response and a rejected response. The accuracy is calculated as the proportion of test samples in which the reward model assigns a higher score to the chosen response than to the rejected one. The detailed results are shown in Table 6. We observe a high accuracy of 83.8% for the reward model, demonstrating its effectiveness in aligning with the intended criteria. Compared to the strong reward models in Table 5, our model achieves the highest accuracy, highlighting the advantage of leveraging RAG-specific reward data.
Furthermore, we observe a consistent accuracy across the 3 tasks, indicating that the reward model could jointly learn the preference signal from diverse tasks and domains. Notably, the reward model achieves the highest accuracy on the Data-to-Text task, while its performance is relatively lower on the Summarization task. This difference suggests that comparing structured data with text data is easier for the reward model, while evaluating the quality of a summarization poses a greater challenge. This observation aligns with our intuition and expectations
@@ 6.2 Preference Alignment
WebGLM 74.8 65.2
Yelp 66.4 68.6
Xsum 72.8 65.4
Average 71.3 66.4
@@
WebGLM 66.8 65.8
Yelp 54.4 57.8
XSum 68.2 72.6
Average 63.1 66.2
Table 8: The RLHF results on the 3 tasks. The win rate is calculated as the proportion of test samples where the GPT-4o prefers the response generated by the post-trained policy model.
@@
To measure the improvement brought by alignment training, we first sample responses from our held-out test set using both the initial models and the post-trained policy models. Thus, for each prompt, we have paired responses from the two models. These pairs are then evaluated using the reward model, and we calculate the proportion of cases where the responses from the post-trained models are preferred. Furthermore, we ask GPT-4o to compare the pairs based on the criteria introduced in Section 3. For both metrics, a baseline of 50% indicates no improvement in the policy models. The calculated scores are referred to as the win rate against the initial model.
The experiment results are shown in Table 7 and Table 8. We observe a clear improvement in the policy models after a single iteration of RAFT. Both the reward model and GPT-4o agree that generations align more closely with the RAG metrics, as the average win rate is significantly above 50%. These results highlight the effectiveness of our dataset and the reward model.
However, we also observe some imbalance in the reward model’s learning compared to GPT-4o. As shown in the tables, the differences in win rates between the two evaluators are minimal for the WebGLM and XSum test sets. However, the difference is more pronounced for the Yelp dataset. This again suggests that the reward model faces greater challenges in evaluating the quality of structured data, such as JSON-format data.
@@ 7 Conclusion
In this paper, we introduce RAG-Reward, a high-quality preference dataset designed for Retrieval-Augmented Generation (RAG). Our dataset is generated through a novel automated AI annotation pipeline, leveraging both open-source and proprietary models to enhance generalization and versatility. To ensure fair and reliable evaluations, we use GPT-4o to assess generation quality based on four key metrics carefully selected by human experts. The dataset spans multiple domains, including Question Answering, Data-to-Text, and Summarization, resulting in a large-scale and diverse benchmark. The experimental results show strong alignment with human evaluations, demonstrating the effectiveness of RAG-Reward in reward modeling and reinforcement learning. These findings highlight the potential of our dataset to advance both the evaluation and generation of RAG systems. To foster further research, we publicly release the dataset to the community.

### v2

@@ 5 Limitation of Existing Reward Models
In this section, we evaluate several existing reward models in our test set. We select models from RewardBench (Lambert et al., 2024) known for their strong performance in assessing aspects such as helpfulness, safety, and reasoning. We examine their performance on diverse RAG scenarios using our curated test set (Table 1), and demonstrate their limitation of the evaluation on RAG domains.
The experiment results are shown in Table 5. While many of the listed reward models achieve accuracy higher than 90% in evaluating chat, safety, and reasoning tasks, their overall accuracy in RAG scenarios is below 80%. This underscores a significant gap between mainstream reward models and the unique requirements of RAG tasks. We also observe inconsistent performance across tasks for the reward models, with better results on Data-to-Text and Summarization tasks (e.g. the Yelp dataset and the XSum dataset) compared to the QA task (WebGLM dataset), suggesting that current reward models are not uniformly capable across different RAG scenarios. Interestingly, several reward models that achieve state-of-the-art (SOTA) performance in reasoning and safety evaluations, as shown on the leaderboard [^2] https://huggingface.co/spaces/allenai/reward-bench, do not perform well on RAG tasks. For example, models like URM-LLaMa-3.1-8B (Lou et al., 2024) and Skywork-Reward-Llama-3.1-8B-v0.2 (Liu et al., 2024) underperform in the RAG domain.
In contrast, UltraRM-13b (Cui et al., 2024), which performs suboptimally in reasoning and safety evaluations but excels in assessing helpfulness and instruction-following, achieves the top accuracy on RAG tasks. This suggests that reward models trained primarily on reasoning tasks may not generalize effectively to evaluating RAG-specific generations. Most of the existing reward models could not excel in expressing the preference in RAG scenarios. Domain-specific training data are therefore essential to address this gap and improve RAG performance evaluation.
@@ 6.1 Reward Modeling
During the test stage, each test sample contains a chosen response and a rejected response. The accuracy is calculated as the proportion of test samples in which the reward model assigns a higher score to the chosen response than to the rejected one. The detailed results are shown in Table 6. We observe a high accuracy of 84.3% for the reward model, demonstrating its effectiveness in aligning with the intended criteria. Compared to the strong reward models in Table 5, our model achieves the highest accuracy, highlighting the advantage of leveraging RAG-specific reward data.
Furthermore, we observe a consistent accuracy across the 3 tasks, indicating that the reward model could jointly learn the preference signal from diverse tasks and domains. Notably, the reward model achieves the highest accuracy on the Data-to-Text task, while its performance is relatively lower on the Question-Answering task. This difference suggests that comparing structured data with text data is easier for the reward model, while evaluating the quality of a long-form QA poses a greater challenge. This observation aligns with our intuition and expectations.
@@ 6.2 Preference Alignment
WebGLM 64.2 62.6
Yelp 77.8 75.0
Xsum 65.6 64.0
Average 69.2 67.2
@@
WebGLM 61.2 60.0
Yelp 80.2 73.6
XSum 63.8 63.2
Average 68.4 65.6
Table 8: The RLHF results on the 3 tasks. The win rate is calculated as the proportion of test samples where the o3-mini prefers the response generated by the post-trained policy model.
@@
To measure the improvement brought by alignment training, we first sample responses from our held-out test set using both the initial models and the post-trained policy models. Thus, for each prompt, we have paired responses from the two models. These pairs are then evaluated using the reward model, and we calculate the proportion of cases where the responses from the post-trained models are preferred. Furthermore, we ask o3-mini to compare the pairs based on the criteria introduced in Section 3. For both metrics, a baseline of 50% indicates no improvement in the policy models. The calculated scores are referred to as the win rate against the initial model.
The experiment results are shown in Table 7 and Table 8. We observe a clear improvement in the policy models after a single iteration of RAFT. Both the reward model and o3-mini agree that generations align more closely with the RAG metrics, as the average win rate is significantly above 50%. These results highlight the effectiveness of our dataset and the reward model. The ratings across the 3 tasks from the reward model is very similar to the ratings from o3-mini, showing that our reward model learns the rationale of rating from o3-mini.
However, we also observe some imbalance in learning for the policy model across tasks. As shown in the tables, there are differences in the win rate across 3 tasks differences. Specifically, the win rate for Yelp could reach 80% while the other 2 are only above 60%, even though they are trained on the same number of samples for each task. The comparison reveals that the difficulties are different across RAG scenarios.
WebGLM Yelp XSum Average
Human 62.0 70.0 66.0 66.0
Table 9: The human evaluation results of the policy model (Mistral) after RAFT on 3 tasks. The agreement is calculated as the proportion of test samples where the generation after RAFT is preferred by humans.
6.3 Human Evaluation
To further validate the improvement of the policy model, we leverage human evaluation of the generations from it. Specifically, we select the model which trained from Mistral-7B-Instruct with one iteration of RAFT, and the Mistral itself as the reference. We adopt the same evaluation strategy introduced in Section 6.2 and replace the reward model and o3-mini with human annotators. Due to the expense, we select 50 samples of each dataset for human labeling. From Table 9, we observe an agreement far above 50%, indicating the effiveness of the RAFT training to improve the policy model on RAG domains. Compared with the results from the reward model and o3-mini, we discover the same trend across 3 tasks, which further shows the alignment of our dataset with humans.
@@ 7 Conclusion
In this paper, we introduce RAG-Reward, a high-quality preference dataset designed for Retrieval-Augmented Generation (RAG). Our dataset is generated through a novel automated AI annotation pipeline, leveraging both open-source and proprietary models to enhance generalization and versatility. To ensure fair and reliable evaluations, we use o3-mini to assess generation quality based on four key metrics carefully selected by human experts. The dataset spans multiple domains, including Question Answering, Data-to-Text, and Summarization, resulting in a large-scale and diverse benchmark. The experimental results show strong alignment with human evaluations, demonstrating the effectiveness of RAG-Reward in reward modeling and reinforcement learning. These findings highlight the potential of our dataset to advance both the evaluation and generation of RAG systems. To foster further research, we will publicly release the dataset to the community.
8 Limitations
In this paper, we constructed a large-scale, high-quality dataset tailored for RAG scenarios and demonstrated the effectiveness of our proposed pipeline through RLHF experiments. However, due to computational constraints, we did not conduct large-scale RLHF training or implement more complex algorithms such as PPO. Future work could explore training larger reward models and incorporating iterative-DPO or PPO to further enhance performance in RAG domains.